A car drove away from its stationary position with a constant acceleration of 5.0 m s-2. 

Plot its motion in the velocity-time graph below for the first 5.0 seconds of travel.

 

(a)   How long does it take the car to reach a speed of

20.0 ms-l?

(b)  How far did the car travel during this time?


1
Expert's answer
2022-05-18T08:32:04-0400

(a)

"t=v/a=20.0/5.00=4.0\:\rm s"

(b)

"d=at^2/2=5.0*4.0^2/2=40\:\rm m"